# Integral cohomology rings of weighted Grassmann orbifolds and rigidity properties

### 11 Mar 2024 v2

@@ Abstract.
In this paper, we introduce ‘Plücker weight vector’ and establish the definition of a weighted Grassmann orbifold $\mbox{Gr}_{\mathbf{b}}(k,n)$, corresponding to a Plücker weight vector ‘$\mathbf{b}$’. We achieve an explicit classification of weighted Grassmann orbifolds up to certain homeomorphism in terms of the Plücker weight vectors. We study the integral cohomology of $\mbox{Gr}_{\bf b}(k,n)$ and provide some sufficient conditions such that the integral cohomology of $\mbox{Gr}_{\bf b}(k,n)$ has no torsion. We describe the integral equivariant cohomology ring of divisive weighted Grassmann orbifolds and compute all the equivariant structure constants with integer coefficients. Eminently, we compute the integral cohomology rings of divisive weighted Grassmann orbifolds explicitly.
@@
Plücker weight vector, weighted Grassmann orbifold, $q$-CW complex, Plücker permutation, lens space, integral cohomology, integral equivariant cohomology, equivariant Schubert basis, structure constant
@@ 1. Introduction
Grassmann manifolds are the central objects of study in algebraic topology, algebraic geometry, and enumerative geometry, see [12, 14, 15]. Consider the $n$-dimensional complex vector space $\mathbb{C}^{n}$ and a positive integer $k$ satisfying $1\leq k<n$. The space comprising all $k$-dimensional subspaces in the complex $n$-dimensional space $\mathbb{C}^{n}$ possesses the structure of a smooth manifold. This manifold is known as a complex Grassmann manifold (in short, Grassmann manifold in this paper) and is denoted by $\mbox{Gr}(k,n)$. Grassmann manifold $\mbox{Gr}(k,n)$ can be equivariantly embedded into the complex projective space $\mathbb{C}P^{m}$, where $m+1={n\choose k}$. This embedding is known as the Plücker embedding, presenting $\mbox{Gr}(k,n)$ as a projective algebraic variety. The image of the Grassmann manifold in the projective space is described by a set of homogeneous polynomial equations known as Plücker relations. The non-zero points in $\mathbb{C}^{m+1}$ that satisfy Plücker relations are referred to as Plücker coordinates.
@@
In this article, we delve into the concept of Plücker coordinates, which are uniquely defined by a set of homogeneous polynomial equations referred to as Pl”ucker relations, as illustrated in (2.5). This exploration leads us to introduce Plücker weight vector ${\bf{b}}=(b_{0},b_{1},\dots,b_{m})\in(\mathbb{Z}_{\geq 0})^{m+1}$, where $m+1={n\choose k}$ for two positive integers $k$ and $n$ such that $k<n$. Then corresponding to a Plücker weight vector ${\bf{b}}$, we define a topological space $\mbox{Gr}_{\bf{b}}(k,n)$. We derive that the definition of $\mbox{Gr}_{\bf b}(k,n)$ generalizes the definition of weighted Grassmannian in [1] as well as weighted Grassmann orbifold in [7]. We show that $\mbox{Gr}_{\bf b}(k,n)$ has an orbifold structure and call this a weighted Grassmann orbifold. This broadens the class of weighted Grassmann orbifolds in terms of Plücker weight vectors. We introduce Plücker permutation, which produces several invariant $q$-CW complex structures on $\mbox{Gr}_{\bf b}(k,n)$. Moreover, we characterize weighted Grassmann orbifolds up to certain homeomorphism in terms of Plücker weight vectors and Plücker permutations. We discuss the integral cohomology of weighted Grassmann orbifolds. Inspired by the work of [13], one can ask the following.
Question 1.1.
Is the integral cohomology of a weighted Grassmann orbifold concentrated in even degrees without torsion? If this happens, can we calculate the integral cohomology rings of weighted Grassmann orbifolds?
The author and Sarkar [7] gave an affirmative answer to the first part of this question for divisive $\mbox{WGr}(k,n)$ and computed the generalized $T^{n}$-equivariant cohomology ring of divisive $\mbox{WGr}(k,n)$, where $T^{n}$ is the standard compact abelian torus of dimension $n$. Moreover, they found an equivariant Schubert basis for the equivariant cohomology ring $H_{T^{n}}^{*}(\mbox{WGr}(k,n);\mathbb{Z})$ as $H^{*}_{T^{n}}(\{pt\};\mathbb{Z})$ module. In this article, we explicitly calculate the equivariant structure constants with respect to the equivariant Schubert basis in $H_{T^{n}}^{*}(\mbox{Gr}_{\bf b}(k,n);\mathbb{Z})$ with integer coefficients. This paves the way for computing the cup product for the cohomology ring of a divisive weighted Grassmann orbifold with integer coefficients. The paper is organized as follows.
@@
$L^{\prime}(b_{i};{\bf{b}}^{(i)})\hookrightarrow X^{i-1}\to X^{i}$
for the lens complex $L^{\prime}(b_{i};{\bf{b}}^{(i)})$, see Proposition 2.12.
In Section 3, we introduce Plücker permutations $\sigma$ on $\{0,1,\dots,m\}$ those permute the Plücker coordinates. Then we show that all Plücker weight vectors are stable under Plücker permutations. We prove that two weighted Grassmann orbifolds are weakly equivariantly homeomorphic if their Plücker weight vectors differ by a Plücker permutation, see Theorem 3.4. This provides several different $q$-CW complex structures of a weighted Grassmann orbifold. Conversely, we prove the following.
@@
If there exists a homeomorphism $f$ between two weighted Grassmann orbifolds ${\rm{Gr}}_{\bf b}(k,n)$ and ${\rm{Gr}}_{\bf c}(k,n)$ such that the $f$ maps every coordinate element to some coordinate element then ${\bf b}$ and ${\bf c}$ are same up to a scalar multiplication and a permutation.
The coordinate elements are the set of all fixed points for the several group actions on $\mbox{Gr}_{\bf b}(k,n)$. Thus the above Theorem is classifies the weighted Grassmann orbifolds up to weakly equivariant homeomorphism, see Corollary 3.12.
In Section 4, for a prime $p$, we deduce the integral cohomology of any weighted Grassmann orbifold has no $p$-torsion under some hypothesis on the Plücker weighted vector, see Theorem 4.1. Moreover, we prove thae following.
@@
We study some non-trivial examples of $\mbox{Gr}_{\bf{b}}(2,4)$ and $\mbox{Gr}_{\bf{b}}(2,5)$ such that the integral cohomologies of those weighted Grassmann orbifolds have no torsion and concentrated in even degrees. We define divisive weighted Grassmann orbifolds in terms of Plücker permutations and Plücker weight vectors. Then we show that the integral cohomology of divisive Grassmann orbifolds has no torsion and is concentrated in even degrees.
In Section 5, we describe the $T^{n}$-equivariant cohomology ring of divisive $\mbox{Gr}_{\bf{b}}(k,n)$ with integer coefficients as well as a basis $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ for the aforementioned ring as a $\mathbb{Z}[y_{1},\dots,y_{n}]$ module following [7]. Subsequently, we can define the equivariant structure constant ${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ satisfies the following.
If ${\bf{b}}=(1,1,\dots,1)$, the above basis leads to a basis $\{\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ for $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$. We find the relation between $\{\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ and the basis considered by Knutson and Tao in [14] for the equivariant cohomology ring $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$. Initially we compute the equivariant structure constants $\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ with respect to the basis $\{\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$, see Lemma 5.7. Then using Vietoris Begle mapping theorem from [19] we get a basis $\{p\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ for $H_{T^{n+1}}^{*}(Pl(k,n);\mathbb{Q})$ as a $\mathbb{Q}[y_{1},y_{2},\dots,y_{n},z]$ module, where $p\widetilde{\mathbb{S}}_{\lambda^{i}}$ is the isomorphic image of $\widetilde{\mathbb{S}}_{\lambda^{i}}$. Then we deduce that ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}$ is the isomorphic image of $p\widetilde{\mathbb{S}}_{\lambda^{i}}$ after evaluating at $z=0$. We discuss the equivariant Pieri rule in $H_{T^{n}}^{*}(\mbox{Gr}_{\bf b}(k,n);\mathbb{Z})$. Moreover we compute the equivariant structure constants with respect to the basis $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$.
@@
${\bf{b}}\widetilde{\mathscr{C}}_{i~{}j}^{\ell}=\sum_{\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{P:\partial P=\Delta_{\sigma_{r}\lambda^{i}~{}\sigma_{r}\lambda^{j}}^{\sigma_{r}\lambda^{q}}}\sum_{s=0}^{|P|}a^{(P)}_{s}{\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}.$
Here $P$ is a puzzle and and $\delta P$ is the boundary of the the labeled equilateral triangle induced from $P$. The coefficients $a^{(P)}_{s}$ and ${\bf{b}}\widetilde{\mathscr{C}}_{q~{}g^{s}}^{\ell}$ can be computed using the equvariant structure cnstants $\widetilde{\mathscr{C}}_{i~{}j}^{\ell}$ in $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$ and the equivariant Pieri rule in $H_{T^{n}}^{*}(\mbox{Gr}_{\bf b}(k,n);\mathbb{Z})$ respectively, see (5.6) and (5.13). We demonstrate the positivity of equivariant structure constants; see Theorem 5.16. Using the upper triangularity of every basis element ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}$, we prove that:
@@
In Section 6, we work out on a basis $\{{\bf b}S_{\lambda^{i}}\}_{i=0}^{m}$ for the integral cohomology groups of divisive weighted Grassmann orbifolds. Finally we compute the structure constant ${\bf{b}}\mathscr{C}_{i~{}j}^{\ell}$ with respect to the above basis with integer coefficients:
@@
Here $\mathscr{C}_{i~{}j}^{\ell}$ is the structure constant in $H^{*}(\mbox{Gr}(k,n);\mathbb{Z})$ and $\mathscr{D}_{i~{}j}^{\ell~{}q}$ is described in (6.2). This describes the integral cohomology ring of a divisive $\mbox{Gr}_{\bf{b}}(k,n)$ explicitly. In particular, we describe the explicit computation of the structure constants and the integral cohomology ring of a divisive $\mbox{Gr}_{\bf{b}}(2,4)$.
@@ 2. Building sequences of weighted Grassmann orbifolds
In this Section, we introduce the notion of Plücker weight vector, which we use to give a new definition of weighted Grassmann orbifold. Then, we show that this definition includes the weighted Grassmannian discussed in [1] as well as [7]. We discuss a building sequence of $\mbox{Gr}_{\bf{b}}(k,n)$ that arises from a $q$-CW complex structure.
For two positive integers $k$ and $n$ such that $k<n$, we recall a Schubert symbol $\lambda$ is a sequence of $k$ integers $(\lambda_{1},\lambda_{2},\dots,\lambda_{k})$ such that $1\leq\lambda_{1}<\lambda_{2}<\cdots<\lambda_{k}\leq n$. A pair $(s,s^{\prime})$ is called a reversal of $\lambda$ if $s\in\lambda,s^{\prime}\notin\lambda$ and $s>s^{\prime}$. The set of all reversals of $\lambda$ is denoted by $\rm{rev}(\lambda)$. If $(s,s^{\prime})\in\rm{rev}(\lambda)$, then $(s,s^{\prime})\lambda$ denote the Schubert symbol obtained by replacing $s$ by $s^{\prime}$ in $\lambda$ and ordering the later set.
@@
Then the Schubert symbols for $k<n$ form a lattice with respect to this partial order $\preceq$’.
There are ($n\choose k$ many Schubert symbols for two positive integers $k$ and $n$ such that $k<n$. Let $m={n\choose k}-1$. Now consider a total order
@@
on the set of all Schubert symbols which preserve the partial order $\preceq$’ i.e., if $\lambda^{i}\preceq\lambda^{j}$ then $i<j$. For example, one can consider the dictionary order on the set of Schubert symbols. Then, it is a total order that preserves the partial order $\preceq$ defined in (2.1).

### 09 Mar 2026 v3

@@ Abstract.
In this paper, we introduce ‘Plücker weight vector’ and establish the definition of a weighted Grassmann orbifold ${\rm Gr}_{\bf b}(k,n)$, corresponding to a Plücker weight vector $\bf b$}’. We achieve an explicit classification of weighted Grassmann orbifolds up to certain homeomorphism in terms of the Plücker weight vectors. We study the integral cohomology of ${\rm Gr}_{\bf b}(k,n)$ and provide some sufficient conditions such that the integral cohomology of ${\rm Gr}_{\bf b}(k,n)$ has no torsion. We explicitly describe the formula of the equivariant structure constants with respect to the equivariant Schubert basis in equivariant cohomology ring of divisive weighted Grassmann orbifolds with integer coefficients. Eminently, we compute the integral cohomology rings of divisive weighted Grassmann orbifolds explicitly.
@@
Plücker weight vector, weighted Grassmann orbifold, $q$-CW complex, Plücker permutation, integral cohomology, equivariant cohomology, equivariant Schubert basis, structure constant
@@ 1. Introduction
Grassmann manifolds are the central objects of study in algebraic topology, algebraic geometry, and enumerative geometry, see [14, 16, 17]. Consider the $n$-dimensional complex vector space $\mathbb{C}^{n}$ and a positive integer $k$ satisfying $1\leq k<n$. The space comprising all $k$-dimensional subspaces in $\mathbb{C}^{n}$ possesses the structure of a smooth manifold. This manifold is known as a complex Grassmann manifold (in short, Grassmann manifold in this paper) and is denoted by $\mbox{Gr}(k,n)$. Grassmann manifold $\mbox{Gr}(k,n)$ can be equivariantly embedded into the complex projective space $\mathbb{C}P^{m}$, where $m+1={n\choose k}$. This embedding is known as the Plücker embedding, presenting $\mbox{Gr}(k,n)$ as a projective algebraic variety. The image of the Grassmann manifold in the projective space is described by a set of homogeneous polynomial equations known as Plücker relations. The non-zero points in $\mathbb{C}^{m+1}$ that satisfy Plücker relations are referred to as Plücker coordinates.
@@
In this article, we delve into the concept of Plücker coordinates, which are defined as solutions to a system of homogeneous polynomial equations known as Plücker relations, as illustrated in (2.5). This perspective naturally leads us to introduce Plücker weight vector ${\bf{b}}=(b_{0},b_{1},\dots,b_{m})\in(\mathbb{Z}_{\geq 0})^{m+1}$, where $m+1={n\choose k}$ for two positive integers $k$ and $n$ such that $k<n$. Associated to each Plücker weight vector $\bf{b}$}, we define a topological space $\mbox{Gr}_{\bf{b}}(k,n)$. The space $\mbox{Gr}_{\bf b}(k,n)$ has an orbifold structure and call this a weighted Grassmann orbifold. We introduce the notion of Plücker permutation, which yields several invariant $q$-CW complex structures on $\mbox{Gr}_{\bf b}(k,n)$. Moreover, we classify weighted Grassmann orbifolds up to certain homeomorphism in terms of Plücker weight vectors and Plücker permutations.
The main goal of this paper is to compute the integral cohomology ring of weighted Grassmann orbifolds with integer coefficients. The author and Sarkar [7] introduced divisive weighted Grassmann orbifolds and computed the generalized $T^{n}$-equivariant cohomology ring of these spaces with integer coefficients, where $T^{n}$ is the standard compact abelian torus of dimension $n$. There exists an equivariant Schubert basis for the equivariant cohomology ring divisive weighted Grassmann orbifolds as $H^{*}_{T^{n}}(\{pt\};\mathbb{Z})$ module. In this article, we explicitly calculate the equivariant structure constants with respect to the equivariant Schubert basis in $H_{T^{n}}^{*}(\mbox{Gr}_{\bf b}(k,n);\mathbb{Z})$ with integer coefficients. This paves the way for computing the cup product for the cohomology ring of a divisive weighted Grassmann orbifold with integer coefficients. The paper is organized as follows.
@@
$L^{\prime}(b_{i};{\bf{b}}^{(i)})\hookrightarrow X^{i-1}\to X^{i},$
see Proposition 2.12, where $L^{\prime}(b_{i};{\bf{b}}^{(i)})$ is a lens complex defined in [15].
In Section 3, we introduce Plücker permutations, permutations $\sigma$ on $\{0,1,\dots,m\}$ those act on the Plücker coordinates. Then we show that all Plücker weight vectors are invariant under Plücker permutations. Moreover, Plücker permutation provides several different $q$-CW complex structures of a weighted Grassmann orbifold. We prove the following:
@@
Two weighted Grassmann orbifolds are weakly equivariantly homeomorphic if their Plücker weight vectors differ by a Plücker permutation and a scalar multiplication.
Conversely, we formulate a conjecture asserting that weighted Grassmann orbifolds can be classified up to homeomorphism solely in terms of their Plücker weight vectors together with Plücker permutations. In support of this conjecture, we prove Lemma 3.10, Lemma 3.11 and Lemma 3.12. Moreover, we prove the following regidity theorem.
@@
If two weighted Grassmann orbifolds ${\rm{Gr}}_{\bf b}(k,n)$ and ${\rm{Gr}}_{\bf c}(k,n)$ are homeomorphic such that each coordinate element mapped to a coordinate element then $\bf b$} and $\bf c$} are identical up to a scalar multiplication and a permutation.
The coordinate elements are the set of all fixed points for the several group actions on $\mbox{Gr}_{\bf b}(k,n)$. Thus, the above Theorem classifies the weighted Grassmann orbifolds up to certain weakly equivariant homeomorphism, see Corollary 3.14.
In Section 4, we show that for a prime $p$, the integral cohomology of any weighted Grassmann orbifold contains no $p$-torsion under some hypothesis on the Plücker weighted vector, see Theorem 4.1. We study some non-trivial examples of $\mbox{Gr}_{\bf{b}}(2,4)$ and $\mbox{Gr}_{\bf{b}}(2,5)$ such that the integral cohomologies of those weighted Grassmann orbifolds have no torsion and concentrated in even degrees. We define divisive weighted Grassmann orbifolds in terms of Plücker permutations and Plücker weight vectors. Then, we show that the integral cohomologies of divisive Grassmann orbifolds are torsion-free and is concentrated in even degrees.
In Section 5, we describe the $T^{n}$-equivariant cohomology ring of the divisive $\mbox{Gr}_{\bf{b}}(k,n)$ with integer coefficients and the equivariant Schubert basis $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ for the aforementioned ring as a $\mathbb{Z}[y_{1},\dots,y_{n}]$ module. The main achievement of this section is computing the equivariant structure constants ${\bf{b}}\widetilde{\mathscr{C}}_{i~j}^{\ell}$ with respect to the basis $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ satisfies the following.
(1.2) ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{j}}=\sum_{\ell=0}^{m}{\bf{b}}\widetilde{\mathscr{C}}_{i~j}^{\ell}{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{\ell}}.$
If ${\bf{b}}=(1,\dots,1)$, the above basis leads to the Schubert basis $\{\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ for $H_{T^{n}}^{*}(\mbox{Gr}(k,n);\mathbb{Z})$. We first describe the equivariant structure constants $\widetilde{\mathscr{C}}_{i~j}^{\ell}$ with respect to the basis $\{\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ in (5.3) following [16]. Then using Vietoris-Begle mapping theorem from [20] we obtain a basis $\{P\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$ for $H_{T^{n+1}}^{*}(Pl(k,n);\mathbb{Q})$ as a $\mathbb{Q}[y_{1},y_{2},\dots,y_{n}]$ module, Then we deduce that ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}$ can be obtained by some isomorphic image of $P\widetilde{\mathbb{S}}_{\lambda^{i}}$. Eventually, we compute the equivariant structure constants with respect to the basis $\{{\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}\}_{i=0}^{m}$.
@@
${\bf{b}}\widetilde{\mathscr{C}}_{i~j}^{\ell}=\sum_{\lambda^{\ell}\succeq\lambda^{q}\succeq\lambda^{i},\lambda^{j}}\sum_{R}\sum_{s=0}^{|R|}K(i,j,q,R)\mathcal{L}_{s,R}{\bf{b}}\widetilde{\mathscr{C}}_{q~g^{s}}^{\ell}.$
Here, $R$ is a finite collection of elements in $\{1,2,\dots,n-1\}$. The non-negative integers $K(i,j,q,R)$ are known due to [16], while the explicit formulae for the remaining coefficients $\mathcal{L}_{s,R}$ and ${\bf{b}}\widetilde{\mathscr{C}}_{q~g^{s}}^{\ell}$ are explained in (5.6) and (5.9) respectively. As a corollary, We describe the equivariant Pieri rule in $H_{T^{n}}^{*}(\mbox{Gr}_{\bf b}(k,n);\mathbb{Z})$. We demonstrate the positivity of equivariant structure constants; see Theorem 5.14. Using the upper triangularity of every basis element ${\bf{b}}\widetilde{\mathbb{S}}_{\lambda^{i}}$, we prove that:
@@
In Section 6, we work out on a basis $\{{\bf b}\mathbb{S}_{\lambda^{i}}\}_{i=0}^{m}$ for the integral cohomology groups of divisive weighted Grassmann orbifolds. Finally, we compute the structure constant ${\bf{b}}\mathscr{C}_{i~j}^{\ell}$ with respect to the above basis with integer coefficients:
@@
Here $\mathscr{C}_{i~j}^{\ell}$ is the structure constant in $H^{*}(\mbox{Gr}(k,n);\mathbb{Z})$ and $\mathscr{D}_{i~j}^{\ell~q}$ is described in (6.3). This explicitly describes the integral cohomology ring of a divisive $\mbox{Gr}_{\bf{b}}(k,n)$. We discuss an example to describe the explicit computation of the structure constants and the integral cohomology ring of a divisive $\mbox{Gr}_{\bf{b}}(2,4)$.
In appendix, we explore the integral cohomology of $\mbox{Gr}_{\bf b}(2,4)$. The main result of this section is the following:
Theorem F (Theorem 7.3).
@@ 2. Building sequences of weighted Grassmann orbifolds
In this Section, we introduce the notion of the Plücker weight vector, which we use to give a new definition of the weighted Grassmann orbifold. Then, we show that this definition generalizes the weighted Grassmannian discussed in [1] and [7]. We discuss a $q$-CW complex structure of weighted Grassmann orbifold and explore the Lens complexes those arise in the cofibration induced from the $q$-CW complex structure.
For two positive integers $k$ and $n$ such that $k<n$, we recall a Schubert symbol $\lambda$ is a sequence of $k$ integers $(\lambda_{1},\lambda_{2},\dots,\lambda_{k})$ such that $1\leq\lambda_{1}<\lambda_{2}<\cdots<\lambda_{k}\leq n$. A pair $(s,s^{\prime})$ is called a inversion of $\lambda$ if $s\in\lambda,s^{\prime}\notin\lambda$ and $s<s^{\prime}$. The set of all inversions of $\lambda$ is denoted by $\rm{inv}(\lambda)$. If $(s,s^{\prime})\in\rm{inv}(\lambda)$, then $(s,s^{\prime})\lambda$ denote the Schubert symbol obtained by replacing $s$ by $s^{\prime}$ in $\lambda$ and ordering the later set.
@@
The Schubert symbols for $k<n$ form a lattice with respect to this partial order $\preceq$’.
There are $n\choose k$ many Schubert symbols for two positive integers $k$ and $n$ such that $k<n$. Now consider a total order <`<’ on the set of all Schubert symbols which preserve the partial order $\prec$’. In other words, if $\lambda\prec\mu$ then $\lambda<\mu$. For example, one can consider the dictionary order on the set of Schubert symbols. Then, it is a total order that preserves the partial order $\preceq$ defined in (2.1). Thus we can consider a chain
@@
on Schubert symbols for $k<n$, where $m={n\choose k}-1$.